Below is a sample space for a family with 3 children. BGG stands for
the oldest child being a boy, the middle child a girl, and the youngest
a girl. Use the sample space to answer the question:
What is the probability (in simplest form) that the middle child is a
a boy? _____
Sample Space
BBB
BBG
BGB
BGG
GBB
GBG
GGB
GGG

Answer by jim_thompson5910(13679)

There are 8 possible configurations of having 3 children. Out of those 8, there are 4 possible ways to have a middle child that is a boy. So

P(Boy middle child) = 4/8 = 1/2


So the probability is 1/2 (which is 0.5 or 50%)